# Strong convergence of fully discrete finite element schemes for the stochastic semilinear generalized Benjamin-Bona-Mahony equation driven by additive Wiener noise

### v1

@@ Abstract.
In this article, we have analyzed semi-discrete finite element approximation and full discretization of the Stochastic semilinear generalized Benjamin-Bona-Mahony equation in a bounded convex polygonal domain driven by additive Wiener noise. We use the finite element method for spatial discretization and the semi-implicit method for time discretization and derive a strong convergence rate with respect to both parameters (spatial and temporal). Numerical experiments have also been performed to support theoretical bounds.
@@
Stochastic BBM equation, Additive noise, Generalized BBM equation, Finite element Method, Strong convergence
@@ 1. Introduction
We consider the stochastic semilinear generalized Benjamin-Bona-Mahony(BBM) equation
$du-d(\Delta u)-(\Delta u-\Delta^{2}u)dt=f(u)dt+dW,\quad t>0,\,x\in\mathcal{O},$ (1.1)
@@
where $\mathcal{O}\subset\mathbb{R}^{d},\,d=1,2,3$ is a bounded convex polygonal domain and $\{W(t)\}_{t\geq 0}$ is an $L_{2}(\mathcal{O})$-valued Wiener process with covariance operator $Q$ on a filtered probability space $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$. We assume that $f:L_{2}(\mathcal{O})\to L_{2}(\mathcal{O})$ satisfies the global Lipschitz condition.
The Benjamin-Bona-Mahony (BBM) equation was proposed by Benjamin, Bona, and Mahony in 1972 [1] and describes the approximate unidirectional propagation of long waves in certain dispersive systems. The Benjamin-Bona-Mahony equation is a more satisfactory model than the Korteweg-de Vries equation (KdV) [10] with improved analytical properties, particularly in terms of regularity and numerical stability.
To address more complex dispersive phenomena, the generalized BBM equation incorporates higher-order spatial derivatives, enabling the modeling of intricate wave interactions in systems with advanced dispersive or stabilizing effects. These higher-order terms are critical in applications such as optics and material science, where complex boundary effects dominate. The generalized BBM equation further extends the model by including higher-order spatial derivatives, which account for additional dispersive or stabilizing effects and describe complex wave interactions [24, 23, 19].
The introduction of stochastic forcing represents a critical advancement in the study of these equations. Real-world systems, however, are often subject to random perturbations, such as environmental noise or external fluctuations, which deterministic models cannot fully capture. The study of finite element approximation of stochastic partial differential equations has emerged as a powerful approach in numerical analysis. The finite element approximation of stochastic heat, wave, and Schrödinger equations has been extensively studied in the literature; see, for example [20, 21, 11, 12, 4, 2]. The semi-implicit time discretization scheme is popular recently; see, for example[16, 9, 18]. These works provide a foundation for developing a finite element approximation and full discretization for the stochastic generalized BBM equation, which combines the fourth-order derivative, nonlinearity, and stochasticity.
@@
In this paper, we develop a finite element approximation as well as a semi-implicit time discretization for the fourth-order stochastic generalized BBM equation with respect to the spatial variable and time, deriving the rate of convergence, and investigating the well-posedness of its solutions. Due to the involvement of the dissipative term $\Delta^{2}u$, finding the rate of convergence and proving the existence of the solution of the equation (1.1) is quite challenging. To address this problem. We transform the fourth-order parabolic equation into a system of parabolic-elliptic coupled equations of second order by introducing a new variable $v=u-\Delta u$, i.e
$\begin{cases}dv-\Delta vdt=f(u)dt+dW,\\ v=u-\Delta u.\end{cases}$ (1.2)
Yubin Yan [20] and Mihály Kovács [11] have studied the rate of convergence of finite element approximation as well as full discretization for the stochastic parabolic equations. Motivated by their study, we aim to investigate the convergence rate to both parameters(spatial and temporal) of the stochastic generalized BBM equation.
@@ 2. Setup and Main Results
Let $\mathcal{O}\subset\mathbb{R}^{d},\,d=1,2,3$ be a bounded convex polygonal domain with boundary $\partial\mathcal{O}$. We consider the stochastic semilinear generalized BBM equation
$\begin{cases}du-d(\Delta u)-(\Delta u-\Delta^{2}u)dt=f(u)dt+dW\quad\text{ in }(0,\infty)\times\mathcal{O},\\ u=\Delta u=0\hskip 150.79968pt\qquad\text{ on }(0,\infty)\times\partial\mathcal{O},\\ u(0,x)=u_{0}(x)\hskip 162.18062pt\text{ in }\mathcal{O},\end{cases}$ (2.1)
where $u_{0}$ is $\mathcal{F}_{0}$ measurable random variable. To analyze the above system, we introduce the change of variable $v=u-\Delta u$, which converts equation (2.1) into a stochastic coupled parabolic-elliptic system
@@
where $v_{0}=u_{0}-\Delta u_{0}$. We observe that in the above system, the equation in $v$ is a parabolic equation coupled with a nonlinear forcing term $f(u)$ and a stochastic perturbation. The second equation in $u$ is an elliptic equation coupled with a linear forcing term $v$. In the elliptic equation, $u(t)$ plays the role of a parameter. Thus, it is a family of elliptic equations. We note that solving for $v$ with the initial condition $v_{0}$ yields the solution $u$ associated with the initial condition $u_{0}$. Our analysis will focus on $v$ to derive the solution and establish the convergence rate of the finite element approximation as well as semi-implicit time discretization for the equation (2.1). The constant $C>0$ is a generic constant, possibly varying from line to line.
@@
$v(t)=E(t)v_{0}+\int_{0}^{t}E(t-\tau)f(u(\tau))d\tau+\int_{0}^{t}E(t-\tau)dW(\tau),\quad\text{for }t\geq 0.$ (2.4)
Let $V_{h}$ be a finite-dimensional subspace of $H_{0}^{1}(\mathcal{O})$. It is defined in Section 3.2. The finite element approximation of equation (2.4) is to find $v_{h}\in V_{h}$ such that
$dv_{h}(t)+A_{h}v_{h}(t)dt=\mathcal{P}_{h}f(u_{h}(t))dt+\mathcal{P}_{h}dW(t),\quad t>0,\,v_{h}(0)=\mathcal{P}_{h}v_{0}=v_{0,h},$ (2.5)
@@
$u_{h}(t)=(\mathcal{P}_{h}+A_{h})^{-1}v_{h}(t)\quad\text{ for }t\geq 0.$ (2.6)
The discrete operator $A_{h}$ and the projection $\mathcal{P}_{h}$ are defined in Section 3.2. In analogy with (2.4), the mild solution of the discretized equation (2.5) is expressed as $\mathbb{P}$ a.s.
$v_{h}(t)=E_{h}(t)v_{0,h}+\int_{0}^{t}E_{h}(t-\tau)\mathcal{P}_{h}f(u_{h}(\tau))d\tau+\int_{0}^{t}E_{h}(t-\tau)\mathcal{P}_{h}dW(\tau),\quad\text{ for }t\geq 0.$ (2.7)
@@
Let $f$ be a globally Lipschitz function, $Q$ be trace class operator satisfying $\|A^{(\beta-1)/2}Q^{1/2}\|_{HS}<\infty$ for $\beta\geq 0$, and $u_{0}$ is a $H^{2}(\mathcal{O})$-valued $\mathcal{F}_{0}$ measurable random variable. Then there is a unique mild solution $\{u(t)\}_{t\geq 0}$ to the equation (2.1). In particular $\|u(t)\|_{L_{2}(\Omega,L_{2}(\mathcal{O}))}\leq C$ for $0\leq t\leq T$, where $C=C(T,f)$.
@@
Let $Q$ be a trace-class operator, $f$ a globally Lipschitz continuous function, and let $v_{0}\in L_{2}(\Omega,L_{2}(\mathcal{O}))$. Then, for all $0\leq t\leq T$, there exists a constant $C=C(T,f,Q)$ such that
@@
By applying Lemma (2.2) in combination with finite element techniques developed for deterministic parabolic problems (see, for instance, [17, Theorem 3.5] ), as recalled in Section 3.2, we derive the following convergence rate for the stochastic semilinear generalized BBM equation (2.1). In this analysis, we employ the fractional Sobolev space $\dot{H}^{\beta}$ which is defined in Section (3).
@@
Let $\{u(t)\}_{t\geq 0}$ denote the mild solution of equation (2.1), and let $\{u_{h}(t)\}_{t\geq 0}$ be its finite element approximation given by equation (2.6). Suppose that $\|A^{(\beta-1)/2}Q^{1/2}\|_{\mathrm{HS}}<\infty$ for some $\beta\in[0,2)$, and $v_{0}\in L_{2}(\Omega,L_{2}(\mathcal{O}))$. Then there exists a constant $C=C(v_{0},T,Q,f)>0$ such that for $0\leq\beta<2$,
@@
Assume that (ii) holds. Let $u$ be the mild solution of (2.2). Then for $0\leq\gamma<\beta\leq 1$, we have
@@
$V^{n}-V^{n-1}+kA_{h}V^{n}=k\mathcal{P}_{h}f(U^{n-1})+\int_{t_{n-1}}^{t_{n}}\mathcal{P}_{h}dW(\tau),\,n\geq 1,\,V^{0}=\mathcal{P}_{h}v_{0},$ (2.8)
@@
$\begin{split}&V^{n}=R(kA_{h})V^{n-1}+R(kA_{h})k\mathcal{P}_{h}f(U^{n-1})+\int_{t_{n-1}}^{t_{n}}R(kA_{h})\mathcal{P}_{h}dW(\tau),\quad n\geq 1,\,\\ &V^{0}=\mathcal{P}_{h}v_{0}.\end{split}$ (2.10)
By combining the temporal continuity of the solution with its mean-square stability, we derive the mean-square error estimate for the fully discrete approximation of the stochastic generalized BBM equation.
@@ 3. Preliminaries
$W(t)=\sum_{j=1}^{\infty}\gamma_{j}^{1/2}\beta_{j}(t)e_{j},$ (3.1)
@@
$\left\|\int_{0}^{t}\Phi(\tau)\,dW(\tau)\right\|^{2}_{L_{2}(\Omega,U_{2})}=\int_{0}^{t}\|\Phi(\tau)Q^{1/2}\|^{2}_{\mathcal{L}_{2}(U_{1},U_{2})}\,d\tau.$ (3.3)
@@ 3.2. Finite element approximations
$h=\max_{T\in\mathcal{T}_{h}}h_{T},$
where $h_{T}=\text{diam}(T)\text{ and, }T\text{ is triangle in }\mathcal{O}$. Define $V_{h}$ as the space of continuous, piecewise linear functions associated with $\mathcal{T}_{h}$ that vanish on $\partial\mathcal{O}$. Consequently, we have $V_{h}\subset H^{1}_{0}(\mathcal{O})=\dot{H}^{1}.$
[…]
@@
Next, we recall some standard results from deterministic finite element analysis and operator semigroup theory. The operator A-A generates an analytic semigroup $\{E(t)\}_{t\geq 0}$ on $H$, while its discrete counterpart $A_{h}$, obtained by finite element discretization, generates the analytic semigroup $\{E_{h}(t)\}_{t\geq 0}$ on $V_{h}$. For the deterministic parabolic finite element problem, the following error estimate holds (see, e.g., [17, Theorem 3.5]):
$\|(E_{h}(t)\mathcal{P}_{h}-E(t))\psi\|\leq Ch^{\beta}t^{-\beta/2}\|\psi\|,\quad 0\leq\beta\leq 2,\;\psi\in H.$ (3.4)

### v2

# Strong convergence of finite element approximations for a fourth-order stochastic pseudo-parabolic equation with additive noise

@@ Abstract.
In this article, we analyze semi-discrete finite element approximation and full discretization of a fourth-order stochastic pseudo-parabolic equation in a bounded convex polygonal domain driven by additive Wiener noise. We use the finite element method for spatial discretization and the semi-implicit method for temporal discretization, and obtain strong convergence rates with respect to both the spatial and temporal mesh sizes. Numerical experiments are presented to support the theoretical convergence rates.
@@
Stochastic pseudo-parabolic equation, Wiener noise, Finite element Method, Strong convergence
@@ 1. Introduction
Pseudo-parabolic equations describe a variety of physical and engineering phenomena, including heat conduction in materials with memory, flow in porous media, and dispersive wave propagation [3, 17, 10, 2, 4]. Equations involving mixed spatial and temporal derivatives exhibit behavior that differs from classical parabolic models and are commonly referred to as pseudo-parabolic. This class of equations was designated as “pseudo-parabolic” by Showalter and Ting in 1970 [30]. This paper focuses on a fourth-order stochastic pseudo-parabolic equation with additive noise
(1.1) $du-d(\Delta u)-(\Delta u-\Delta^{2}u)dt=f(u)dt+dW,$
@@
where $f$ is a nonlinear source term and $\{W(t)\}_{t\geq 0}$ is an $L_{2}(\mathcal{O})$-valued Wiener process with covariance operator $Q$ on a filtered probability space $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ with respect to the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$.
For the fourth-order nonlinear pseudo-parabolic equation (1.1) without a noise term, global existence and asymptotic behavior of solutions have been extensively investigated [19, 39, 13, 14] and the references therein. In addition to qualitative analysis, numerical approximations of pseudo-parabolic equations have also attracted considerable attention [35, 1, 22, 25, 23]. Various numerical methods, including finite difference schemes, spatial finite element discretizations, and backward Euler time-stepping methods, have been employed to analyze and approximate such models.
In many realistic applications, however, deterministic models are often influenced by random perturbations arising from environmental fluctuations, measurement errors, or unresolved physical mechanisms. This motivates incorporating stochastic effects into pseudo-parabolic equations via Wiener processes, yielding stochastic pseudo-parabolic models. Theoretical investigations concerning the existence, uniqueness, and asymptotic behavior of solutions to stochastic pseudo-parabolic equations have been reported in [24, 18, 7]. Moreover, stochastic fractional pseudo-parabolic equations have also been studied, and results on the existence and uniqueness of solutions can be found in [34, 31].
Despite the growing body of literature on the qualitative analysis of stochastic pseudo-parabolic equations, numerical approximation for such stochastic models remains relatively underdeveloped. In particular, to the best of our knowledge, strong convergence analysis for fully discrete finite element schemes combined with semi-implicit time discretization applied to fourth-order stochastic pseudo-parabolic equations has not yet been established. The presence of mixed spatial and temporal derivatives, together with stochastic forcing, introduces additional analytical challenges compared with classical stochastic parabolic equations.
By contrast, finite element approximations for stochastic heat, wave, and Schrödinger equations have been extensively investigated in the literature; see, for example, [36, 37, 20, 21, 8, 5]. Moreover, semi-implicit time discretization schemes have recently attracted considerable attention due to their stability and efficiency; see, e.g., [29, 16, 33]. These developments provide a solid foundation for constructing and analyzing fully discrete finite element schemes for the fourth-order stochastic pseudo-parabolic equation considered in this work, which involves the interplay of fourth-order spatial derivatives, nonlinear effects, and stochastic perturbations.
@@
In this paper, we develop a finite element approximation and a semi-implicit time discretization for the fourth-order stochastic pseudo-parabolic equation with respect to both spatial and temporal variables, derive the convergence rates, and investigate the well-posedness of its solutions. Due to the involvement of the dissipative term $\Delta^{2}u$ and mixed derivative, finding the rate of convergence and proving the existence of the solution of the equation (1.1) is quite challenging. To address this problem, we transform the fourth-order stochastic pseudo-parabolic equation into a stochastic coupled parabolic–elliptic system by introducing a new variable $v=u-\Delta u$, i.e
(1.2) $\begin{cases}dv-\Delta vdt=f(u)dt+dW,\\ v=u-\Delta u.\end{cases}$
Main contributions. The main contributions of this paper are summarized as follows.
  • We establish the existence and uniqueness of mild solutions for the fourth-order stochastic pseudo-parabolic equation driven by additive Wiener noise and derive suitable regularity estimates that serve as the foundation for the numerical analysis.
  • We analyze both the semi-discrete finite element approximation in space and the fully discrete scheme that combines a finite element spatial discretization with a semi-implicit time-stepping method. Strong convergence rates are derived with respect to the spatial and temporal mesh sizes.
  • Numerical experiments with explicit examples are presented to validate the theoretical convergence results
@@
Throughout the paper, the constant $C$ denotes a generic positive constant independent of the discretization parameters. Its value may vary from line to line.
@@ 2. Setup and Main Results
Let $\mathcal{O}\subset\mathbb{R}^{d},\,d=1,2,3$ be a bounded convex polygonal domain with boundary $\partial\mathcal{O}$. We consider the fourth-order stochastic pseudo-parabolic equation
(2.1) $\begin{cases}du-d(\Delta u)-(\Delta u-\Delta^{2}u)dt=f(u)dt+dW\quad\text{ in }(0,\infty)\times\mathcal{O},\\ u=\Delta u=0\hskip 150.79968pt\qquad\text{ on }(0,\infty)\times\partial\mathcal{O},\\ u(0,x)=u_{0}(x)\hskip 162.18062pt\text{ in }\mathcal{O},\end{cases}$
where $f:L_{2}(\mathcal{O})\to L_{2}(\mathcal{O})$ is a globally Lipschitz function and $u_{0}$ is an $\mathcal{F}_{0}$ measurable random variable. To analyze the above system, we introduce the change of variable $v=u-\Delta u$, which converts equation (2.1) into a stochastic coupled parabolic-elliptic system
@@
where $v_{0}=u_{0}-\Delta u_{0}$. We observe that in the above system, the equation in $v$ is a parabolic equation coupled with a nonlinear forcing term $f(u)$ and a stochastic perturbation. The second equation in $u$ is an elliptic equation coupled with a linear forcing term $v$. In the elliptic equation, $u(t)$ plays the role of a parameter. Thus, it is a family of elliptic equations. We note that solving for $v$ with the initial condition $v_{0}$ yields the solution $u$ associated with the initial condition $u_{0}$. Our analysis will focus on $v$ to derive the solution and establish the convergence rate of the finite element approximation as well as semi-implicit time discretization for the equation (2.1).
@@
(2.4) $v(t)=E(t)v_{0}+\int_{0}^{t}E(t-\tau)f(u(\tau))d\tau+\int_{0}^{t}E(t-\tau)dW(\tau),\quad\text{for }t\geq 0.$
Let $V_{h}$ be a finite-dimensional subspace of $H_{0}^{1}(\mathcal{O})$. It is defined in Subsection 3.2. The finite element approximation of equation (2.4) is to find $v_{h}\in V_{h}$ such that
(2.5) $dv_{h}(t)+A_{h}v_{h}(t)dt=\mathcal{P}_{h}f(u_{h}(t))dt+\mathcal{P}_{h}dW(t),\quad t>0,\,v_{h}(0)=\mathcal{P}_{h}v_{0}=v_{0,h},$
@@
(2.6) $u_{h}(t)=(\mathcal{P}_{h}+A_{h})^{-1}v_{h}(t)\quad\text{ for }t\geq 0.$
The discrete operator $A_{h}$ and the projection $\mathcal{P}_{h}$ are defined in Subsection 3.2. In analogy with (2.4), the mild solution of the discretized equation (2.5) is expressed as $\mathbb{P}$ a.s.
(2.7) $v_{h}(t)=E_{h}(t)v_{0,h}+\int_{0}^{t}E_{h}(t-\tau)\mathcal{P}_{h}f(u_{h}(\tau))d\tau+\int_{0}^{t}E_{h}(t-\tau)\mathcal{P}_{h}dW(\tau),\quad\text{ for }t\geq 0.$
@@
Let $f$ be a globally Lipschitz function, $Q$ be trace class operator satisfying $\|A^{(\beta-1)/2}Q^{1/2}\|_{HS}<\infty$ for $\beta\geq 0$, and $u_{0}$ is an $\mathcal{F}_{0}$ measurable $H^{2}(\mathcal{O})$-valued random variable. Then there is a unique mild solution $\{u(t)\}_{t\geq 0}$ to the equation (2.1). In particular $\|u(t)\|_{L_{2}(\Omega,L_{2}(\mathcal{O}))}\leq C$ for $0\leq t\leq T$, where $C=C(T,f)$.
@@
Let $Q$ be a trace-class operator, $f$ a globally Lipschitz continuous function, and let $v_{0}\in L_{2}(\Omega,L_{2}(\mathcal{O}))$. Then, for $0\leq t\leq T$, there exists a constant $C=C(T,f,Q)$ such that
@@
By applying Lemma 2.2 in combination with finite element techniques developed for deterministic parabolic problems (see, for instance, [32, Theorem 3.5] ), as recalled in Subsection 3.2, we derive the following convergence rate for the fourth-order stochastic pseudo-parabolic equation (2.1). In this analysis, we employ the fractional Sobolev space $\dot{H}^{\beta}$, defined in Section 3.
@@
Let $\{u(t)\}_{t\geq 0}$ denote the mild solution of equation (2.1), and let $\{u_{h}(t)\}_{t\geq 0}$ be its finite element approximation given by equation (2.6). Suppose that $\|A^{(\beta-1)/2}Q^{1/2}\|_{\mathrm{HS}}<\infty$ for some $\beta\in[0,2)$, and $v_{0}\in L_{2}(\Omega,\dot{H}^{\beta})$. Then there exists a constant $C=C(v_{0},T,Q,f)>0$ such that for $0\leq\beta<2$,
@@
Assume that the growth condition (ii) holds. Let $u$ be the mild solution of (2.2). Then for $0\leq\gamma<\beta\leq 1$, we have
@@
(2.8) $V^{n}-V^{n-1}+kA_{h}V^{n}=k\mathcal{P}_{h}f(U^{n-1})+\int_{t_{n-1}}^{t_{n}}\mathcal{P}_{h}dW(\tau),\,n\geq 1,\,V^{0}=\mathcal{P}_{h}v_{0},$
@@
(2.10) $\begin{split}&V^{n}=R(kA_{h})V^{n-1}+R(kA_{h})k\mathcal{P}_{h}f(U^{n-1})+\int_{t_{n-1}}^{t_{n}}R(kA_{h})\mathcal{P}_{h}dW(\tau),\quad n\geq 1,\,\\ &V^{0}=\mathcal{P}_{h}v_{0}.\end{split}$
By combining the temporal continuity of the solution with its mean-square stability, we derive the mean-square error estimate for the fully discrete approximation of the fourth-order stochastic pseudo-parabolic equation.
@@ 3. Preliminaries
(3.1) $W(t)=\sum_{j=1}^{\infty}\gamma_{j}^{1/2}\beta_{j}(t)e_{j},$
@@
(3.3) $\left\|\int_{0}^{t}\Phi(\tau)\,dW(\tau)\right\|^{2}_{L_{2}(\Omega,U_{2})}=\int_{0}^{t}\|\Phi(\tau)Q^{1/2}\|^{2}_{\mathcal{L}_{2}(U_{1},U_{2})}\,d\tau.$
@@ 3.2. Finite element approximations
$h=\max_{K\in\mathcal{T}_{h}}h_{K},$
where $h_{K}=\text{diam}(K)\text{ and, }K\text{ is triangle in }\mathcal{O}$. Define $V_{h}$ as the space of continuous, piecewise linear functions associated with $\mathcal{T}_{h}$ that vanish on $\partial\mathcal{O}$. Consequently, we have $V_{h}\subset H^{1}_{0}(\mathcal{O})=\dot{H}^{1}.$
Since $\mathcal{O}$ is assumed to be convex and polygonal, the triangulations can be precisely aligned with $\partial\mathcal{O}$. Moreover, this ensures the elliptic regularity property $\|w\|_{H^{2}(\mathcal{O})}\leq C\|Aw\|$ for $w\in D({A})$, see [15]. We now recall fundamental results from the finite element theory (see [9, 6]). The norms are denoted as $\|\cdot\|_{s}=\|\cdot\|_{\dot{H}^{s}}$.
@@
Next, we recall some standard results from deterministic finite element analysis and operator semigroup theory. The operator A-A generates an analytic semigroup $\{E(t)\}_{t\geq 0}$ on $H$, while its discrete counterpart $A_{h}$, obtained by finite element discretization, generates the analytic semigroup $\{E_{h}(t)\}_{t\geq 0}$ on $V_{h}$. For the deterministic parabolic finite element problem, the following error estimate holds (see, e.g., [32, Theorem 3.5]):
(3.4) $\|(E_{h}(t)\mathcal{P}_{h}-E(t))\psi\|\leq Ch^{\beta}t^{-\beta/2}\|\psi\|,\quad 0\leq\beta\leq 2,\;\psi\in H.$